The proportion of people in a given community who have a certain disease is 0.005. A test

is available to diagnose the disease. If a person has the disease, the probability that the test

will produce a positive signal is 0.99. If a person does not have the disease, the probability

that the test will produce a positive signal is 0.01. If a person tests positive, what is the

probability that the person actually has the disease?


1
Expert's answer
2022-01-13T10:12:35-0500

Solution:

Consider N be the persons which have true diseases and + sign indicates the positive sign of the test. And consider "N^{\prime}" be the persons which have not true diseases and + sign indicates the positive sign of the test. So, the conditional probability that if a person tests positive (X) given actually disease is calculated as:

"\begin{gathered}

P(X \mid+)=\frac{P(X \cap+)}{P(+)} \\

=\frac{P(+X) P(X)}{P(+X) P(X)+P\left(+X^{\prime}\right) P\left(X^{\prime}\right)} \\

=\frac{0.99 \times 0.005}{(0.99 \times 0.005)+(0.01 \times(1-0.005))} \\

=0.3322

\end{gathered}"

The required conditional probability that if a person test positive given actual disease is approximately 0.3322 .